OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. x2 + x +10 = 0 A. {3.2} B. {-5.5, -2} C. {-3.2, 3.2} D. no real number solution

hero (hero):

\[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

hero (hero):

But actually, just do b^2 - 4ac to make sure that it is even necessary to do quad form

hero (hero):

(1)^2 - 4(1)(10) = 1 - 40 = -39 Since discriminant is negative, this means that there are no real solutions